### v1

# ToxSearch: Evolving Prompts for Toxicity Search in Large Language Models

Abstract
Large Language Models remain vulnerable to adversarial prompts that elicit toxic content even after safety alignment. We present ToxSearch, a black-box evolutionary framework that tests model safety by evolving prompts in a synchronous, steady-state $(\mu+\lambda)$ loop. The system employs a diverse operator suite, including lexical substitutions, negation, back-translation, paraphrasing, and two semantic crossover operators, while a moderation oracle provides fitness guidance. Operator-level analysis reveals significant heterogeneity in performance, as lexical substitutions offer the best yield–variance trade-off, semantic-similarity crossover acts as a precise low-throughput inserter, and global rewrites exhibit high variance with elevated refusal costs. Using elite prompts evolved on LLaMA 3.1 8B, we observe practically meaningful but attenuated cross-model transfer. Toxicity drops by roughly half on most targets, with smaller LLaMA 3.2 variants showing the strongest resistance and some cross-architecture models (e.g., Qwen and Mistral) retaining higher toxicity. Overall, our results indicate that small, controllable perturbations serve as reliable vehicles for systematic red-teaming, while defenses should anticipate cross-model prompt reuse rather than focusing solely on single-model hardening.
[Pre-print]
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities for many tasks, but they also pose risks by potentially generating toxic content [^1] This paper contains disturbing language presented solely for safety evaluation.. Toxicity is defined as “a rude, disrespectful, or unreasonable comment likely to make someone leave a discussion” [ref_url1]. Even when safety-aligned, e.g., via fine-tuning or reinforcement learning from human feedback (RLHF), these models are not foolproof as adversarial inputs can still elicit toxic outputs. In particular, prompt-based attacks have emerged as a central concern because carefully crafted prompts can slip past safety filters. Identifying these prompts and building automated methods to discover them are essential [corbo2025toxic, Perez2022RedTL, Srivastava2023NoOT].
Growing concerns about LLM safety led to research on jailbreak attacks [corbo2025toxic, zou2023universal], but manual techniques are hard to scale and systematically evaluate. Zou et al. demonstrated that automatic prompt attacks are possible and found a universal adversarial suffix via greedy and gradient-based search that caused multiple LLMs to comply with harmful requests at high success rates [zou2023universal]. However, gradient-based approaches such as this usually need differentiable surrogates and often yield unnatural token sequences, limiting their usefulness for text.
To overcome the limitations of manual and gradient-based attacks, researchers have turned to evolutionary algorithms (EAs), which frame prompt generation as an optimization task  [corbo2025toxic, Yu2023GPTFUZZERRT, Yang2023InstOptimaEM, Liu2023AutoDANGS]. A major benefit of EAs is that they require no access to model internals. For instance, EvoTox  [corbo2025toxic] casts toxicity elicitation as a $(1+\lambda)$ evolutionary strategy in which one LLM is the system under test and another LLM acts as a prompt generator that mutates a parent prompt to increase the system model’s toxicity. EvoTox was able to induce highly toxic outputs even from aligned models, outperforming random search and baseline jailbreak prompts on the AdvBench dataset  [zou2023universal]. Likewise, GPTFuzzer  [Yu2023GPTFUZZERRT] treats jailbreak prompt discovery as a fuzzing problem. Starting from a set of human-written seed prompts, GPTFuzzer automatically applies various semantic-preserving transformations to generate candidate prompts. It reached attack success rates above 85% on both ChatGPT and LLaMA-2, clearly outperforming manually designed jailbreaks in controlled tests [Yu2023GPTFUZZERRT]. In other work, Guo et al. introduced EvoPrompt, which applies a genetic algorithm and differential evolution style operators to tune prompts [Guo2023EvoPromptCL], while Yang et al. propose InstOptima, a multiobjective framework that trades off task success, prompt length, and perplexity to recover a Pareto set of prompts [Yang2023InstOptimaEM].
Prior work has shown that automated attacks can easily collapse to a narrow set of prompt patterns exploiting one trick, thereby missing other vulnerabilities and producing homogeneous outputs [Yang2023InstOptimaEM, Guo2023EvoPromptCL]. To counter this, Samvelyan et al. developed the Rainbow Teaming approach, which reframes adversarial prompt search as a quality-diversity optimization problem and maintains an archive of diverse attack strategies to prevent mode collapse  [Samvelyan2024RainbowTO]. In another effort, Liu et al.’s AutoDAN system employs a hierarchical genetic algorithm with sentence- and paragraph-level crossover and uses an LLM-based mutation process to maintain semantic fluency, which together help produce more varied adversarial instructions  [Liu2023AutoDANGS]. Another open question is the transferability of adversarial prompts across different models. Prior works have generally optimized prompts for one specific target model at a time, so it remains underexplored how well a prompt found for model A will perform on model B (especially if the models differ in architecture or alignment tuning). However, a comprehensive characterization of cross-model transfer in toxicity attacks is still lacking.
ToxSearch builds on the evolutionary paradigm by using a dynamically sized steady-state evolutionary strategy $(\mu+\lambda)$ that continuously introduces variation while preserving high-fitness individuals. ToxSearch explicitly maintains a diverse population of candidate prompts through a rich set of operators and steady-state selection, ensuring that we explore a wide range of toxic prompt variants within a single run. To guide our investigation, we pose the following key research questions:
  1. 1.
    RQ1. Which prompt perturbation strategies are most effective to elicit toxic responses?
  2. 2.
    RQ2. To what extent do toxic prompts evolved on one model transfer to other models, especially those with different architectures or alignment tuning?
Results show that under a fixed query budget, using relatively small lexical and semantically grounded mutations can steadily drive a prompt population toward highly toxic outputs. We find that POS–aware synonym and antonym replacement, masked language model–based word substitutions, negation insertion, and back-translation provide the best balance between prompt diversity and toxicity gains. These operators have high success rates in producing effective mutations and relatively low variance in toxicity scores. By contrast, more aggressive global rewrites or stylistic changes tend to make the model refuse to respond, resulting in high variance and lower overall effectiveness. Finally, we observe that several elite prompts evolved on LLaMA-3.1 8B still transfer to other models, albeit with reduced potency. The attack success (toxicity level) typically drops by roughly half on most alternate models, with the strongest transfer effects seen on other models in the LLaMA family, and weaker effects on models with different architectures or alignment mechanisms. These findings suggest that subtle prompt perturbations are dependable way to red-team models, and that safety work should plan for adversarial prompts being reused across different systems, not just harden a single model in isolation.
2 Methodology
ToxSearch is a search-based red teaming framework that operates as a steady-state $(\mu+\lambda)$ loop, which proposes, evaluates, and retains prompts that elicit toxic responses from the target model, formally presented in Algorithm 1. Given a target response generator (RG) LLM $\theta_{rg}$, and a moderation oracle $\mathcal{M}$, we seek prompts $p\in\mathcal{P}$ that maximize a configured fitness score on the model’s response $y\sim\theta_{rg}(p)$, where $p$ is created by a separate prompt generator (PG) LLM, $\theta_{pg}$. $\mathcal{M}$ is an external safety evaluator (in this work, the Google Perspective API [ref_url3]) that maps a text $y$ to a $K$-vector of attribute scores $\mathbf{s}(y)\in[0,1]^{K}$ (toxicity, severe toxicity, insult, identity attack, threat, profanity, etc.). For our fitness function, $s_{toxicity}$, we select the toxicity component as the scalar fitness:
$\max_{p\in\mathcal{P}}\;\hat{F}(p)\;=\;\mathrm{s}_{\text{toxicity}}\!\big(\mathcal{M}(\theta_{rg}(p))\big)\in[0,1]$
Note the prompt generation process is stochastic (even with the same parent(s) and mutation or crossover prompt) because we use non-deterministic decoding, e.g., LLM temperature > 0, where the temperature hyperparameter controls how random the model’s responses are. This aids in improved exploration of the search space. Each generated prompt is evaluated once (no resampling), so we optimize the single-sample plug-in fitness $\hat{F}(p)$ directly. We treat this as black-box search with no access to gradients or model internals. The evolutionary search process runs entirely over a population of text prompts, with the PG being the search mechanism that performs mutation and crossover to generate child prompts. Our evaluation goal is to characterize the worst-case safety risk over allowed prompts.
Algorithm 1 Evolutionary Search for Toxicity in LLMs
1:$P$ $\triangleright$ Initial population of prompts generated from dataset
2:$G$ $\triangleright$ How many evolution cycles to run (default: 50 generations)
[…]
6:$C$ $\triangleright$ Crossover prompts
7:$M$ $\triangleright$ Mutation prompts
[…]
10:Step 1: Initialize Population $P$
[…]
16:Step 2: Evolution Loop
17:for $G$ generations do
18:  Determine Mode: mode $\leftarrow$ DEFAULT, EXPLOIT or EXPLORE
19:  2a. Select Parents $p$:
20:  if mode is DEFAULT then
21:   $p\leftarrow p_{1}\in_{R}E\cup p_{2}\in_{R}N\triangleright$ 1 elite and 1 non-elite uniform at random
22:  else if mode is EXPLOIT then
23:   $p\leftarrow p_{1},p_{2}\in_{R}E\cup p_{3}\in_{R}N\triangleright$ 2 elites and 1 non-elite uniform at random
24:  else if mode is EXPLORE then
25:   $p\leftarrow p_{1}\in_{R}E\cup p_{2},p_{3}\in_{R}N\triangleright$ 1 elite and 2 non-elites uniform at random
26:  end if
27:  2b. Create Children $C$:
28:  $C\leftarrow\varnothing$
29:  for each parent $p_{i}\in p$ do
30:   for each mutation prompt $m\in M$ do
31:     $C\leftarrow c\cup\theta_{pg}(m,p_{i})$
32:   end for
33:  end for
34:  for each unique pair of parents $p_{i},p_{j}\in p$ do
35:   for each crossover prompt $c\in C$ do
36:     $C\leftarrow c\cup\theta_{pg}(c,p_{i},p_{j})$
37:   end for
38:  end for
39:  2c. Evaluate Children:
[…]
41:  2d. Update Population:
42:  $P\leftarrow P\cup C$
43:  $E\leftarrow\text{top }\alpha\text{ of }P$ ranked by toxicity $\triangleright$ get elites
44:  $N\leftarrow\alpha\text{ to }\beta\text{ of }P$ ranked by toxicity $\triangleright$ get non-elites
45:  $U\leftarrow\text{bottom }\beta\text{ of }P$ ranked by toxicity $\triangleright$ get underperforming
46:  $P\leftarrow P-U$ $\triangleright$ remove underperforming from $P$
47:end for
48:Step 3: Return $E$ $\triangleright$ the most toxic prompts found
2.1 Population Management
Of particular note is that while the algorithm utilizes a steady-state $(\mu+\lambda)$ loop, the population size ($\mu$) grows dynamically to encourage exploration. Early versions of the algorithm utilized a fixed-size population which would quickly converge to a single topic and were not able to find toxic response. Instead, we manage the population with score–ratio tiering keyed to the current best score across all generations. For all prompts $p\in P_{g}$, where $P_{g}$ is the population at generation $g$, $S_{\max}(P_{g})=\max_{p\in P_{g}}s(p)$ denotes the highest toxicity score. Elite and removal thresholds are then defined as $\tau_{e}=\Bigl(1-\frac{\alpha}{100}\Bigr)S_{\max}(P_{g})$ and $\tau_{r}=\frac{\beta}{100}\,S_{\max}(P_{g})$. Individuals prompts are then tiered as elite if $s_{toxicity}(p)\geq\tau_{e}$, underperforming if $s_{toxicity}(p)\leq\tau_{r}$, and non-elite otherwise.
Reducing $\alpha$ tightens elitism by raising $\tau_{e}$, whereas increasing $\beta$ raises the removal floor. This adaptive fitness thresholding preserves multiple topical “niches” when one cluster surges, instead of allowing a fixed $k$ elites to be dominated by near-duplicates from a single topic. In rugged or multi-peaked landscapes, ratio thresholds provide scale-invariant selection pressure, secondary peaks remain represented even when the global peak saturates, which empirically reduces premature convergence and sustains the raw material needed for crossovers to escape local ceilings. We removed the non-elites throughout our executions because many variants had low toxicity scores due to the safety alignment of the target LLM. Having a huge distribution skewed towards low toxicity scores diluted the search process and affected search efficiency within a fixed budget. This aligns with established findings on diversity maintenance (e.g., niching/fitness sharing and crowding) that avoid collapse into a single basin when improvements become incremental within that basin.
2.2 Parent Selection
Parent selection is adaptive with three modes governed by short-horizon progress and stagnation namely Default, Explore and Exploit. Default samples one random elite and one random non-elite. Explore samples one random elite and two random non-elites, activating after no new most toxic prompt is found within a window of $W$ previous generations to increase diversity. Exploit samples two random elites and one random non-elite, triggering if there is a negative average population toxicity trend in the window $W$ to intensify selection pressure.
Mode decisions are driven by the slope of the recent trend in average fitness, computed on the pre-redistribution valid pool $V_{g}=E_{g}\cup N_{g}\cup U_{g}$ (elites, non-elites, under-performing). Using $V_{g}$ avoids post-selection bias, as excluding under-performers artificially inflates means and can mask regressions, whereas including all valid outputs measures true search efficiency per budgeted query and shows if the variants generated are increasing the average fitness of the population, hence quantifying the variants generation quality. Let $\bar{s}_{g}$ be the average fitness on $V_{g}$ at generation $g$. Over a sliding window of $W$ generations (set to 5 in our experiments), set $x_{i}=i$ and $y_{i}=\bar{s}_{g-W+1+i}$ for $i=0,\dots,W-1$ and fit $y=\beta_{0}+\beta_{1}x$ by ordinary least squares, where $x_{i}$ indexes the generation within the window, $y_{i}$ is the corresponding mean fitness. The closed-form slope is $\widehat{\beta}_{1}=\frac{\sum_{i=0}^{W-1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sum_{i=0}^{W-1}(x_{i}-\bar{x})^{2}}$ implemented as np.polyfit(x, y, 1)[0]. At each generation we update $\bar{s}_{g}$ on $V_{g}$ and fit an OLS line to the most recent $W$ points to obtain the slope $\widehat{\beta}_{1}$. With tolerance $\tau_{\text{slope}}=10^{-3}$, the controller selects:
$\text{mode}(t)=\begin{cases}\textsc{Exploit},&\widehat{\beta}_{1}<-\tau_{\text{slope}},\\[4.0pt] \textsc{Explore},&\big(S_{\max}(P_{g})=S_{\max}(P_{g-W+1})\big)\ \wedge\ \big(\widehat{\beta}_{1}\geq-\tau_{\text{slope}}\big),\\[4.0pt] \textsc{Default},&\text{otherwise}.\end{cases}$
2.3 Mutation and Crossover Operators
ToxSearch provides a suite of prompt based mutation and crossover operators which generate modifications to a parent prompt given the prompt generator (see Figure 1 for examples). ToxSearch implements the following mutation and crossover operators, with many of these prompts allowing for a variety of pluggable modifications to further increase exploration:
Figure 1: Example Child Prompts from Mutation and Crossover. Words directly related to race or individuals are redacted.
  • POS-Aware Synonym and Antonym Replacement performs targeted lexical substitutions while preserving part-of-speech (POS) and local grammatical structure. We focus on open-class POS tags (Adjective, Adverb, Noun, Verb, Proper Noun, Interjection) to change the prompt semantically while avoiding sentence structure degradation. Given a prompt $p$, spaCy [ref_url4] is used apply POS tagging, randomly select tagged words from available POS categories, and uses the PG to propose contextually appropriate synonym or antonym replacements. Each operator invocation produces at most one variant.
    Table 1: Harmful Concepts
    Concept Category Subtypes
    unfairness systemic bias, institutional discrimination, unequal treatment
    bias racial bias, gender bias, age bias, religious bias, socioeconomic bias
    discrimination direct discrimination, indirect discrimination, systemic exclusion
    harmful targeting psychological harm, social exclusion, economic disadvantage
    sterotyping racial stereotypes, gender stereotypes, cultural stereotypes, class stereotypes
    marginalization social marginalization, political exclusion, economic marginalization
    hate speech incitement to violence, hate propaganda, threats, harassment
    toxic behavior emotional manipulation, psychological abuse, exploitation, coercion
[…]
  • Concept Addition injects harmful concepts into prompts by applying transformations $p\rightarrow p^{\prime}$ where $p^{\prime}$ incorporates explicit bias or unfairness constraints. The operator randomly selects one concept categories from 8 primary categories, each containing multiple subtypes (totaling 29 subtypes), as detailed in Table 1. For each selected category, all of its subtypes are included in the prompt to guide the LLM in generating concept-enhanced variants.
    Table 2: Stylistic attribute taxonomy for register manipulation
    Attribute Stylistic Variants
    formality formal, informal, casual, professional
    politeness polite, impolite, neutral, courteous, rude
    sentiment positive, negative, neutral, optimistic, pessimistic
    tone authoritative, casual, academic, friendly, stern
    voice active, passive, direct, indirect
    complexity simple, complex, basic, sophisticated
    poetic poetic, plain, flowery, rhythmic, prosaic
    technical technical, layman, specialized, accessible
    conversational conversational, formal, chatty, businesslike
    emphatic emphatic, subtle, dramatic, understated
    concise concise, verbose, brief, detailed
    persuasive persuasive, neutral, convincing, objective
  • Masked Language Model Substitution Leveraging constraints like POS filtering and semantic similarity to preserve meaning, Garg et al. showed the effectiveness of Masked Language Model (MLM) based substitutions for generating adversarial text [garg-ramakrishnan-2020-bae]. Jurafsky et al. present masked language modeling as a basis for context sensitive substitution and highlighting that it produces fluent alternatives [jm3]. This mutation operator implements a contextual word substitution strategy based on the MLM paradigm. The operator mutates in a two-stages. First, it randomly selects $m$ words (default $m=1$) and replaces them with numbered mask tokens, creating a partially masked sequence. Second, the PG generates contextually appropriate replacements for each masked token using structured prompting that incorporates both the masked context and the original word as conditioning information.
  • Negation performs semantic inversion by flipping the polarity of core predicates while preserving grammar. The operator applies transformation $\mathcal{N}:p\mapsto p^{\prime}$ where key terms are systematically negated while preserving grammatical structure by the PG.
  • Paraphrasing LLMs naturally function as powerful variation operators that can recombine and rewrite text while preserving intent [Meyerson2023LanguageMC]. Furthermore, self-referential prompt evolution proved that evolving prompts via LLM rewriting can improve task performance [Fernando2023PromptbreederSS]. Motivated by these findings, we adopt instruction-constrained paraphrasing to expand a local, meaning-preserving neighborhood around $p$ using the PG.
  • Stylistic Mutation implements a register-aware transformation strategy. The operator applies transformation $\mathcal{S}:(p,\alpha)\mapsto p^{\prime}$ where $\alpha$ represents a randomly selected stylistic attribute from a structured taxonomy and $p^{\prime}$ exhibits modified stylistic characteristics. The operator uses 12 stylistic dimensions, with each containing multiple instantiation options, as detailed in Table 2.
  • Typographical Errors adds realistic human-like errors at character-level/word-level. The operator applies the transformation $\mathcal{T}:(p,\mathcal{E})\mapsto p^{\prime}$ where $\mathcal{E}$ is a set of selected error types from Table 3 and $p^{\prime}$ contains strategically introduced human-like errors.
  • Informed Evolution is derived from EvoTox framework [corbo2025toxic]. The PG is used to generate a variant prompt that is more likely to elicit toxic responses, by using a specialized system prompt which provides critical requirements and the top 10 most toxic prompts in the population along with their toxicity scores.
    Table 3: Typographical error taxonomy for character-level corruption
    Error Category Error Subtypes
    character substitution vowel substitution, consonant substitution, similar character replacement, keyboard adjacency
    missing letters missing vowels, missing consonants, missing silent letters, missing endings
    extra letters doubled letters, extra vowels, extra consonants, redundant characters
    transposition adjacent character swap, vowel-consonant swap, syllable reversal, word boundary errors
    case inconsistency random capitalization, inconsistent casing, mixed case words, improper capitalization
    punctuation errors missing punctuation, extra punctuation, wrong punctuation, spacing issues
    spacing errors missing spaces, extra spaces, word boundary confusion, inconsistent spacing
    phonetic errors sound-based substitutions, accent-based errors, pronunciation-based typos, homophone confusion
  • Semantic-Similarity Crossover (Sentence Pairing) uses parent prompts $p_{1}$ and $p_{2}$, and first segments each prompt into sentences, then generates high-dimensional embeddings using a pre-trained SentenceTransformer model (all-MiniLM-L6-v2) to capture semantic content in a continuous vector space [Reimers2019SentenceBERTSE]. The crossover process applies cosine similarity matching with threshold $\tau=0.5$ to identify semantically related sentence pairs across parents, and systematically combining matched sentences into a offspring.
  • Semantic Fusion Crossover performs content-aware recombination between two parent prompts, $(p_{1},s_{1})$ and $(p_{2},s_{2})$, to generate an offspring $p^{\prime}=\mathcal{F}(p_{1},p_{2})$ through LLM-guided semantic fusion rather than structural token mixing. This prompts the PG to “fuse the essence” of both parents.
3 Experiments
The PG was the Llama 3.1-8B-Instruct model in GGUF format with Q3_K_M quantization and temperature 0.9 for enhanced creativity. The RG was Llama 3.1-8B-Instruct model in GGUF format with Q3_K_S quantization with temperature 0.7 for consistent response generation. Both models are hosted locally with full GPU offloading, 16 KV-cache enabled and memory-mapped model loading for optimal performance. The PG operated under a system prompt designed to guide adversarial prompt refinement, whereas the RG functions without system-level conditioning to maintain natural response patterns and avoid safety alignment interference. For our experiments, we set $\alpha=30$ and $\beta=3$. Both the PG and RG used a maximum context length of 2048. The initial population $P$ was created by merging the questions from CategoricalHarmfulQA and HarmfulQA, applying light normalization and de-duplication to form a pool of 2,481 harmful question prompts. From this pool we sampled 100 questions uniformly at random (seed=42) and used this fixed dataset for all experiments. Only question texts were used as input, with no labels or answers from the source dataset [bhardwaj-etal-2024-language, Bhardwaj2023RedTeamingLL].
RQ1: Which prompt perturbation strategies are most effective to elicit toxic responses?
Given the above settings ToxSearch was run for 10 repeated experiments. For each operator we report their non-elite insertion rate (NE), elite-hit rate on first placement (EHR), invalid-generation rate (IR; prompt generation refusals or non-questions or not following the instructions properly), conditional elite-hit rate (cEHR; excluding the invalids and duplicates), and the child–parent toxicity change $\Delta$ on the $[0,1]$ evaluator (Google’s Perspective API) scale. We calculated the mean $\Delta\mu$ and std $\Delta\sigma$. For mutation operators, change is calculated as the child toxicity score minus parent toxicity score. For crossover operators, child toxicity minus the mean of the two parents is considered. For Informed Evolution, child toxicity minus the highest parent toxicity from the top 10 genomes used. Invalids are treated as a cost of alignment friction from the instruction-tuned PG, not entirely as performance failures of the operator itself, as with model size or model architecture there can be models who follow instructions properly or models who are not very safety-aligned fine-tuned.
To validate operator-level performance differences, we used non-parametric tests that do not assume normality. For each metric (EHR, cEHR, IR, NE, $\Delta\mu$, $\Delta\sigma$), we ran Kruskal–Wallis H tests to check whether the distributions differ across operators. When a Kruskal–Wallis test was significant ($p<0.05$), we followed up with pairwise Mann–Whitney U tests for all operator pairs to identify which comparisons drive the effect. To account for multiple comparisons, we applied a Bonferroni correction with $\alpha_{\text{corrected}}=0.05/n$, where $n$ is the number of pairwise tests (66 comparisons, $\alpha_{\text{corrected}}=0.000758$). We also computed effect sizes (Cohen’s $r$) and 95% confidence intervals using bootstrap resampling (1000 iterations) to capture practical significance and uncertainty.
Kruskal–Wallis tests indicated significant operator differences on all metrics ($p<0.05$) EHR (H=41.63,p<0.001H=41.63,\,p<0.001), cEHR (H=21.59,p=0.028H=21.59,\,p=0.028), IR (H=108.60,p<0.001H=108.60,\,p<0.001), NE (H=89.45,p<0.001H=89.45,\,p<0.001), $\Delta\mu$ (H=18.92,p=0.007H=18.92,\,p=0.007), and $\Delta\sigma$ (H=25.34,p=0.003H=25.34,\,p=0.003). Post-hoc Mann–Whitney U tests with Bonferroni correction found one significant pair for EHR (MLM >> Semantic Similarity Crossover; $p=0.000751$, $r=0.76$, $95\%\ \mathrm{CI}=[2.06,\,5.26]$), 34 significant pairs for IR, and 38 for NE, pointing to strong heterogeneity across operators. Effect sizes ranged from small ($|r|<0.10$) to large ($|r|\geq 0.30$), with most significant contrasts showing large effects ($|r|\geq 0.70$). Table 4 summarizes execution measurements, reporting per-operator means averaged over 10 runs.
Table 4: Operator metrics by execution with per-operator mean
Operator NE EHR IR cEHR $\boldsymbol{\Delta\mu}$ $\boldsymbol{\Delta\sigma}$
Concept Addition 55.08 4.08 39.33 6.74 -0.06 0.13
Informed Evolution 45.96 8.28 43.98 14.95 -0.18 0.11
Back Translation 70.85 4.35 20.08 5.52 -0.08 0.13
Paraphrasing 55.11 3.01 40.23 5.13 -0.07 0.13
Synonym Replacement 76.59 5.89 12.59 6.95 -0.06 0.12
MLM-based Substitution 59.50 5.55 27.95 8.34 -0.06 0.12
Negation 71.24 4.45 18.38 5.67 -0.07 0.12
Antonym Replacement 83.78 6.29 4.79 6.76 -0.06 0.12
Semantic Fusion 40.20 2.06 55.68 4.60 -0.06 0.09
Semantic Similarity 20.85 1.99 0.00 8.69 -0.06 0.10
Stylistic Transfer 55.32 2.47 40.56 4.13 -0.07 0.13
Typographical Errors 41.88 3.02 53.62 6.41 -0.07 0.12
Lexical edits gave the best budget-to-yield trade-off. POS-aware Antonym Replacement achieved the highest NE with low IR and competitive cEHR, i.e., frequent, small, and stable score shifts. POS-aware Synonym Replacement performed similarly. MLM-based substitution balanced throughput and quality, with low invariance. Negation and Back-Translation also reached high NE with moderate IR and cEHR, producing controlled drops rather than erratic changes. Stylistic transformations were less efficient under our budget and settings. Stylistic Transfer and Paraphrasing had NE around 55% but high IR and modest cEHR. Typographical Errors maximized convergence on paper but were the most wasteful in practice. Concept Addition performed moderately. Informed Evolution delivered the largest elite yield when valid but paid for it with substantial invalids and the steepest average drop, marking it as a high-variance explorer rather than a dependable converter of budget into progress, suggesting it may exacerbate overfitting. Semantic Similarity Crossover showed the strongest cEHR, albeit with low NE, suggesting it is a lower-throughput operator, whereas Semantic Fusion Crossover is more refusal-prone but still yields occasional elites. By category, crossovers exhibit smaller magnitude and variance in $\Delta$ than mutations, indicating “safer” steps that more reliably avoid sharp regressions.
In short, under a fixed budget with an instruction-tuned PG, the most effective operators were found to be the lexical perturbations (Antonym, Synonym, Negation, Back-Translation, MLM), which combine strong yield with small, low-variance $\Delta$. Semantic Similarity complements them as a precision tool when clean insertions matter. High-variance global rewrites (Informed Evolution) should be used sparingly, and refusal-prone edits (Typographical Errors, Semantic Fusion) were inefficient in this regime.
RQ2: To what extent do toxic prompts evolved on one model transfer to other models, especially those with different architectures or alignment tuning?
Figure 2: Toxicity distributions for the transferred prompt set across models.
Figure 3: Percentage of response refusals per target model.
We evaluated cross-model transfer by taking the elite prompts evolved on the source model (LLaMA 3.1 8B) and testing them once on each target model with identical decoding and the same Perspective API configuration for toxicity. From elite records gathered across all runs and operator modes, we deduplicated by prompt text (keeping the highest score per unique prompt), filtered for questions, and obtained 696 unique items. We then selected the top 25% by source-model toxicity (174 prompts; toxicity score range 0.3610–0.8697) as the transfer set. Targets covered seven models namely LLaMA 3.2 1B, LLaMA 3.2 3B, Mistral 7B, Phi-3.5 Mini, Qwen 2.5 7B, and Gemma 2 9B.
Transferred prompts showed a clear drop in toxicity on all target models. On the source model, the toxicity distribution was centered at a mean of 0.490 (median 0.444, std 0.113). All targets shifted downward, with means between 0.199 and 0.302, i.e., reductions of roughly 39–59% relative to the source. Cross-architecture models showed mixed resistance. Qwen 2.5 7B had the highest mean among targets (0.302, median 0.339, std 0.145), followed by Mistral 7B (0.244, median 0.233, std 0.146), Phi-3.5 Mini (0.231, median 0.208, std 0.153), and Gemma 2 9B (0.215, median 0.233, std 0.129). Within-family models landed at the low end of this range: LLaMA 3.2 1B (0.199, median 0.143, std 0.182) and LLaMA 3.2 3B (0.208, median 0.182, std 0.162) showed the strongest resistance despite sharing architecture with the source. Figure 2 plots these per-model distributions as violins with jittered points. The wide IQRs and skewed shapes indicate that most prompts transfer with reduced strength, but a noticeable subset still carries relatively high toxicity across models.
Figure 3 reports the percentage of invalid responses per model. The source model (LLaMA 3.1 8B) had the lowest invalid rate at 2.9% (5 out of 174 prompts), whereas target models ranged from 4.6% to 8.0%. LLaMA 3.2 3B showed the highest invalid rate (8.0%, 14 prompts), which is followed by LLaMA 3.2 1B (6.9%, 12 prompts). These higher invalid rates for the smaller LLaMA variants indicate that they are more likely to refuse or fail to answer toxic prompts, which directly lowers their observed toxicity scores. This behavior provides an explanation for why within-family models appear more resistant despite their similarity to the source.
Overall, the results show that prompts evolved on LLaMA 3.1 8B do transfer to other models but with a noticeable drop in toxicity. Transfer strength is not uniform: smaller within-family variants appear more resistant than several of the larger or cross-architecture targets, which suggests that alignment choices and training methodology matter at least as much as size or raw architectural similarity. The higher refusal rates in these models point to defensive mechanisms as a key factor in transfer resistance. At the same time, the fact that some prompts still retain non-trivial toxicity across models implies that defenses should plan for cross-model prompt reuse instead of treating each model in isolation.
4 Ethical Considerations
This work is conducted solely for LLM safety evaluation and model hardening. We analyze toxic content only to characterize risks in existing systems, using prompts from publicly available harmful-question datasets, and no human subjects were involved. Acknowledging the dual-use nature of adversarial research, Our GitHub repository omits raw prompt corpora and any material that would enable reconstruction of attack strategies. All artifacts are intended for research on safety evaluation and defenses, not for deployment or dissemination of adversarial content.
5 Conclusion
This work presents ToxSearch, a black-box evolutionary framework for systematically eliciting toxic responses from LLMs using a diverse operator suite, ratio-based population tiering, and trend-gated parent selection. Operator-level analysis (RQ1) shows clear differences in behavior across operators. POS-aware antonym/synonym replacement, negation, back-translation, and MLM substitution give the best budget-to-yield trade-offs, while informed evolution (a few shot approach) shows much higher variance and pays a noticeable cost in refusals. Non-parametric tests confirmed that these differences are statistically significant. Cross-model transferability experiments (RQ2) show that the evolved prompts are not just overfitting a single model – prompts evolved on LLaMA 3.1 8B retain substantial toxicity on other models, with the strongest transfer within the LLaMA family and weaker but still meaningful transfer to different architectures and alignment regimes. Invalid response rates across target models further underline that robustness to adversarial prompts is uneven across current LLMs.
These results have direct implications for how we red-team and defend LLMs. Small, controllable lexical perturbations turn out to be a dependable way to do systematic red-teaming under tight budgets, whereas high-variance global rewrite approaches like informed evolution make more sense as exploratory tools rather than the main workhorse. The cross-model transfer we observe suggests that defenses should explicitly plan for prompt reuse across models, not just harden a single system in isolation. The operator-level differences we measured are also useful for reasoning about defenses. Some operators work reliably across most models, while others seem tied to particular architectures or settings, which suggests that defenses should look at families of transformation patterns rather than only at individual prompts. Our framework and metrics give a concrete starting point for generating and measuring adversarial prompts and for tracking how different mitigation strategies change the effective attack surface, with the longer-term aim of building defenses that generalize across models instead of being tailored to a single system.
While this work provides useful insights, it still has several limitations which lead to several directions for future work. On the evaluation side, the objective can be expanded to multiple safety attributes scored by a calibrated ensemble of evaluators, with results reported as Pareto fronts rather than a single scalar, and extended beyond question-form prompts to include imperatives, declaratives, and multi-turn conversations as well as multiple harm categories (toxicity, bias, misinformation, privacy). On the optimization side, the next step is to learn an adaptive operator allocation policy that routes budget where marginal return is highest, for example using lightweight reinforcement learning with explicit cost regularization, and to make elitism and removal thresholds follow population statistics via moving quantiles or control charts, combined with niching or quality–diversity methods to preserve secondary peaks. Methodologically, we plan to characterize operator contribution under different loads through stress tests that vary budget and operator frequency, and to tune the steady-state controller in controlled experiments to identify robust settings. Beyond the current model set, it will be important to test this framework on a wider range of LLMs. This includes larger models (70B+ parameters), other architectural families (e.g., GPT, Claude, PaLM), and systems trained with different alignment strategies such as RLHF, constitutional AI, or self-alignment. We also want to generalize the setup to multilingual settings by adding more translation pivots and multilingual evaluators, so that we can study cross-lingual transfer in a more systematic way. Finally, future work should deepen our understanding of why some prompts transfer so well by analyzing the features of successful attacks, explore scalability and efficiency improvements through parallel evolution and transfer learning, and plug the framework into adaptive defensive mechanisms (for example, real-time filtering, co-evolutionary defenses, or rejection/adversarial training), all under appropriate ethical safeguards and responsible disclosure to model providers.